OpenStudy (anonymous):

k^2+5k+6=(k+2)(k+[]) A.3 B.6 C.12 D.4

OpenStudy (anonymous):

what are you trying to find

Nnesha (nnesha):

since the leading coefficient is one find two numbers when u multiply them you should get product of AC(a=leading coefficient , c=constant) and when add/subtract them you should get the middle term

Directrix (directrix):

k^2+5k+6 = (k+2)(k+[]) (k + 2) (k + 3) = (k+2) (k+[]) The (k + 2) factors divide out. (k + 3) = (k+[ ? ]) @biggirl99